# Constraining a de Broglie–Bohm quantum bounce cosmology with Planck data

### v2

@@ Abstract
This work investigates a singularity-free early Universe within the paradigm of quantum cosmology. We develop a bouncing model where the singularity is resolved via the de Broglie–Bohm interpretation of quantum mechanics, which provides a deterministic trajectory for the scale factor through a quantum bounce. The primordial power spectrum for scalar perturbations is derived, incorporating a characteristic modulation (distortion function) imprinted by the nonstandard quantum gravitational dynamics near the bounce. We confront this model with the Planck 2018 cosmic microwave background data, establishing its strong compatibility with observations. Our analysis places a stringent upper bound on the fundamental scale of the bounce $k_{B}$, constraining the parameter space of such quantum cosmological scenarios. Furthermore, the model’s specific scale-dependent anticorrelation between the spectral index and amplitude of perturbations offers a potential mechanism for mitigating the $H_{0}$-$\sigma_{8}$ tension, presenting a testable signature for future cosmological surveys.
@@ I Introduction
Modern cosmology is fundamentally grounded in the hot big bang model, which describes a dynamic, expanding Universe that emerged from an ultrahot and dense primordial state. This framework enjoys significant success, providing a robust explanation for fundamental observations, including the cosmic microwave background (CMB) radiation, the abundance of light elements from big bang nucleosynthesis, and the large-scale structure (LSS) of the Universe. Nevertheless, this model is inherently incomplete, as it extrapolates back to an initial spacetime singularity—a point where the predictions of general relativity break down and energy densities diverge. This singularity represents not merely a mathematical artifact but the fundamental boundary of the theory’s validity, necessitating a paradigm shift to a quantum theory of gravity for a complete description of the origin of the Universe.
@@
A cornerstone of modern cosmology is the presence of a primordial epoch of accelerated expansion, known as cosmic inflation. This paradigm elegantly resolves key fine-tuning problems of the standard big bang model—namely, the horizon, flatness, and magnetic monopole problems [1, 2, 3]. Furthermore, it provides a powerful mechanism for generating the spectrum of primordial quantum fluctuations that seed the large-scale structure observed in the Universe today. Despite its remarkable successes, the standard inflationary framework is typically constructed on a classical spacetime geometry, implicitly assuming its validity immediately after the Planck epoch. Consequently, it does not address the singularity problem and remains divorced from a first-principles quantum gravitational description of the Universe’s very origin.
@@
Bouncing models [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] provide a compelling alternative to cosmic inflation by resolving the initial singularity problem while preserving its observational successes. Quantum cosmology emerges as a natural framework for such models, as the classical description of spacetime breaks down at the high energies where the bounce occurs, necessitating a quantum gravitational treatment. In this picture, the initial singularity is replaced by a nonsingular quantum bounce, forming a bridge between a preceding contracting phase and our observed expanding Universe.
@@
Furthermore, while bouncing models can, in principle, be considered alternatives to inflation, they are not mutually exclusive. A particularly compelling and widely studied scenario is a hybrid approach in which a quantum bounce is followed by a period of inflation. This synthesis addresses the initial singularity problem through the bounce while simultaneously leveraging inflation’s proven successes, such as generating the large-scale structure of the Universe. Among the most rigorously developed approaches in this vein is loop quantum cosmology (LQC), which applies loop quantum gravity techniques to homogeneous spacetimes [19, 20, 21]. It is important to note that a primary application of LQC is not merely to provide a stand-alone bounce, but to study the conditions for and dynamics of a subsequent inflationary epoch. Consequently, LQC is very commonly investigated within the context of this ”bounce-plus-inflation” paradigm, where the quantum-gravitational bounce provides natural initial conditions for the inflaton field [22, 23, 24].
@@
In LQC, the repulsive character of quantum geometry at near-Planckian densities naturally generates a bounce, thus avoiding the singularity. Crucially, this quantum-gravitational transition is not merely a theoretical artifact, but can imprint specific signatures on primordial perturbations. As demonstrated, for example, in Refs. [25, 26, 27, 28, 29], these may include a suppression of power spectrum at large angular scales in the CMB and characteristic modifications to the primordial gravitational wave spectrum, rendering the paradigm possibly testable through future CMB observations.
@@
Complementary to these geometrically motivated approaches like LQC, another framework seeks to resolve the singularity through the direct quantization of the Universe’s wave function, governed by the Wheeler-DeWitt (WDW) equation. The fundamental interpretative challenges of the WDW equation—its timeless nature and the problem of quantum measurement for a closed Universe—motivate the adoption of alternative quantum interpretations. The de Broglie–Bohm (dBB) pilot wave theory [30, 31] is particularly well suited for this task. The dBB formulation of quantum cosmology provides a deterministic description of quantum evolution via well-defined trajectories, guided by the phase of the universal wave function, where the Universe is described by a single, well-defined geometric configuration guided by the wave function through the quantum potential. Even when the wave function is highly nonclassical, the background evolution follows a unique Bohmian trajectory, ensuring a definite spacetime geometry. As a result, the background geometry is never in a superposition, which guarantees the consistency of treating perturbations as quantum fields propagating on a definite spacetime. This property allows cosmological perturbations to be consistently quantized and evolved on an effective semiclassical background, in close analogy with other semiclassical treatments in quantum cosmology [32]. Crucially, the dBB interpretation furnishes an objective, observer-independent description of quantum dynamics without invoking the collapse postulate, thereby offering a coherent ontological framework for quantum cosmology and a potential mechanism for a nonsingular bounce [15, 33].
@@
Building upon this dBB framework, in Ref. [34] a concrete cosmological model featuring a stiff-matter dominated quantum bounce that transitions smoothly into a standard inflationary phase was developed. The model, which considers a flat Friedmann-Lemaître-Robertson-Walker (FLRW) Universe with a scalar field, yields complete analytical solutions spanning the entire history from quantum contraction through the bounce to the end of inflation. This analytical tractability provides a powerful tool for investigating how fundamental quantum parameters of the prebounce wave function—such as the characteristic bounce density and wave packet dispersion—determine the number of inflationary e-folds and thereby set the initial conditions for the formation of large-scale structure.
@@ II.2 Quantum model: de Broglie–Bohm interpretation
which leads, upon imposing the Hamiltonian constraint, to the WDW equation in the kination regime,
@@ III Primordial Scalar Perturbations
The primary challenge is to evolve quantum vacuum fluctuations through the nonsingular bounce and into the subsequent inflationary epoch. The preinflationary dynamics near the bounce imprint a characteristic signature on the perturbation modes that persists throughout the subsequent evolution. To model this process, we segment the history into three distinct dynamical regimes where the background evolution can be well approximated, thus permitting a tractable analytical treatment for the perturbations:
@@ III.1 Bounce phase
where $\alpha$ and $\mathcal{V}_{0}$ are free parameters fixed in such a way that Eq. (20) can become a good approximation to Eq. (LABEL:VT). Since we can analytically solve the modes equation (15) for a function like (20), we expect this to provide a good approximation for the solution in the bounce phase.
@@ III.2 Transition phase
During the transition phase, the Universe smoothly transits from an equation of state with $\omega=1$ (kination) to $\omega=-1$ (inflation). Hence, we can consider $k^{2}\gg a^{\prime\prime}(\eta)/a(\eta)$ in Eq. (15) and the solution for the modes $\mu_{k}$ is simply plane-wave-like,
@@ III.3 Inflationary phase
where $\alpha_{k}$ and $\beta_{k}$ are arbitrary constants, and $H^{\rm(1)}_{\nu}(z)$ and $H^{\rm(2)}_{\nu}(z)$ are the Hankel functions of the first and second kinds, respectively.
@@ III.4 Matching the solutions
Let us now determine all the arbitrary coefficients that appear in the solutions obtained above in each of the three phases. We are, in particular, interested in the coefficients $\alpha_{k}$ and $\beta_{k}$ of the slow-roll inflationary modes, Eq. (26), resulting from the whole evolution beginning far before the bounce, in the contracting phase. All the coefficients are obtained by matching the solutions and are explained below.
@@
We now match the bounce mode solution in the far past with the one in the next phase, the transition phase. This is done by computing the bounce modes, Eq. (III.1), for $\eta-\eta_{\rm B}\gg 0$. To do this, we first note that in the limit $\eta-\eta_{\rm B}\gg 0$,
@@
Hence, the bounce mode result Eq. (III.1) for $\eta-\eta_{\rm B}\gg 0$ can be simplified and we find
@@ IV Analysis and Results
The constraints obtained for the standard cosmological parameters that exhibit a Gaussian-like posterior profile are presented in Table 1, where we also report the total matter density ($\Omega_{m}$) and the amplitude of matter fluctuation ($\sigma_{8}$) parameters. The corresponding 2D confidence levels (CLs) and one-dimensional marginalized posterior distributions are plotted in Fig. 4.
@@
It should be noted that our cosmological analysis also sheds light on the well-known $H_{0}$-$\sigma_{8}$ correlation and its potential breaking within a QC model. As detailed in Table 1, for the $c=0.003$ case, our results indicate a $H_{0}$ value that remains consistent with those found for the $c=0.03$ and $c=0.3$ scenarios. Simultaneously, we observe a slightly lower $\sigma_{8}$ value compared to the higher $c$ values. This particular shift, exhibiting a tendency toward an anticorrelation between $H_{0}$ and $\sigma_{8}$, holds significant relevance in the context of current cosmological tensions [48].
@@ V Conclusions
In this work, we have investigated the preinflationary Universe in the context of quantum cosmology using the de Broglie–Bohm interpretation. Our primary goal was to analyze the evolution of primordial perturbations during a quantum bounce phase and to constrain the model parameters through the observational implications of this theoretical framework.
@@
Finally, we have shown a possible connection of our QC model with the $H_{0}$ and $\sigma_{8}$ tensions. By driving $H_{0}$ toward values more consistent with local measurements, while concurrently achieving a relatively lower $\sigma_{8}$ (compared to other parameter choices within our model), our QC model offers a promising avenue to reconcile these long-standing discrepancies between early- and late-Universe observations of cosmic evolution. This aligns with recent discussions in the literature, where analogous mechanisms in inflationary models have been explored to break such correlations and mitigate cosmological tensions, as discussed, for example, in Ref. [53].
@@ Acknowledgements.
M.B. acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), iniziativa specifica QGSKY. We thank the use of the CAMB code [35] and the Cobaya MCMC sampler [36] in this work. R.O.R. is partially supported by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021.

### v3

# Constraining a de Broglie-Bohm quantum bounce cosmology with Planck data

@@ Abstract
This work investigates a singularity-free early Universe within the paradigm of quantum cosmology. We develop a bouncing model where the singularity is resolved via the de Broglie-Bohm interpretation of quantum mechanics, which provides a deterministic trajectory for the scale factor through a quantum bounce. The primordial power spectrum for scalar perturbations is derived, incorporating a characteristic modulation (distortion function) imprinted by the non-standard quantum gravitational dynamics near the bounce. We confront this model with the Planck 2018 cosmic microwave background data, establishing its strong compatibility with observations. Our analysis places a stringent upper bound on the fundamental scale of the bounce $k_{B}$, constraining the parameter space of such quantum cosmological scenarios. Furthermore, the model’s specific scale-dependent anti-correlation between the spectral index and amplitude of perturbations offers a potential mechanism for mitigating the $H_{0}$-$\sigma_{8}$ tension, presenting a testable signature for future cosmological surveys.
@@ I Introduction
Modern cosmology is fundamentally grounded in the hot Big Bang model, which describes a dynamic, expanding Universe that emerged from an ultra-hot and dense primordial state. This framework enjoys significant success, providing a robust explanation for fundamental observations, including the cosmic microwave background (CMB) radiation, the abundance of light elements from Big Bang nucleosynthesis (BBN), and the large-scale structure (LSS) of the Universe. Nevertheless, this model is inherently incomplete, as it extrapolates back to an initial spacetime singularity — a point where the predictions of general relativity break down and energy densities diverge. This singularity represents not merely a mathematical artifact but the fundamental boundary of the theory’s validity, necessitating a paradigm shift to a quantum theory of gravity for a complete description of the origin of the Universe.
@@
A cornerstone of modern cosmology is the presence of a primordial epoch of accelerated expansion, known as cosmic inflation. This paradigm elegantly resolves key fine-tuning problems of the standard Big Bang model — namely the horizon, flatness, and magnetic monopole problems [2, 3, 4]. Furthermore, it provides a powerful mechanism for generating the spectrum of primordial quantum fluctuations that seed the large-scale structure observed in the Universe today. Despite its remarkable successes, the standard inflationary framework is typically constructed on a classical spacetime geometry, implicitly assuming its validity immediately after the Planck epoch. Consequently, it does not address the singularity problem and remains divorced from a first-principles quantum gravitational description of the Universe’s very origin.
@@
Bouncing models [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] provide a compelling alternative to cosmic inflation by resolving the initial singularity problem while preserving its observational successes. Quantum cosmology emerges as a natural framework for such models, as the classical description of spacetime breaks down at the high energies where the bounce occurs, necessitating a quantum gravitational treatment. In this picture, the initial singularity is replaced by a non-singular quantum bounce, forming a bridge between a preceding contracting phase and our observed expanding universe.
@@
Furthermore, while bouncing models can, in principle, be considered alternatives to inflation, they are not mutually exclusive. A particularly compelling and widely studied scenario is a hybrid approach in which a quantum bounce is followed by a period of inflation. This synthesis addresses the initial singularity problem through the bounce while simultaneously leveraging inflation’s proven successes, such as generating the large-scale structure of the universe. Among the most rigorously developed approaches in this vein is Loop Quantum Cosmology (LQC), which applies loop quantum gravity techniques to homogeneous spacetimes [20, 21, 22]. It is important to note that a primary application of LQC is not merely to provide a standalone bounce, but to study the conditions for and dynamics of a subsequent inflationary epoch. Consequently, LQC is very commonly investigated within the context of this ”bounce-plus-inflation” paradigm, where the quantum-gravitational bounce provides natural initial conditions for the inflaton field [23, 24, 25].
@@
In LQC, the repulsive character of quantum geometry at near-Planckian densities naturally generates a bounce, thus avoiding the singularity. Crucially, this quantum gravitational transition is not merely a theoretical artifact, but can imprint specific signatures on primordial perturbations. As demonstrated, for example, in Refs. [26, 27, 28, 29, 30], these may include a suppression of power spectrum at large angular scales in the CMB and characteristic modifications to the primordial gravitational wave spectrum, rendering the paradigm possibly testable through future CMB observations.
@@
Complementary to these geometrically motivated approaches like LQC, another framework seeks to resolve the singularity through the direct quantization of the Universe’s wave function, governed by the Wheeler-DeWitt (WDW) equation. The fundamental interpretative challenges of the WDW equation — its timeless nature and the problem of quantum measurement for a closed Universe — motivate the adoption of alternative quantum interpretations. The de Broglie-Bohm (dBB) pilot-wave theory [31, 32] is particularly well suited for this task. The dBB formulation of quantum cosmology provides a deterministic description of quantum evolution via well-defined trajectories, guided by the phase of the universal wave function, where the Universe is described by a single, well-defined geometric configuration guided by the wave function through the quantum potential. Even when the wave function is highly nonclassical, the background evolution follows a unique Bohmian trajectory, ensuring a definite spacetime geometry. As a result, the background geometry is never in a superposition, which guarantees the consistency of treating perturbations as quantum fields propagating on a definite spacetime. This property allows cosmological perturbations to be consistently quantized and evolved on an effective semiclassical background, in close analogy with other semiclassical treatments in quantum cosmology [33]. Crucially, the dBB interpretation furnishes an objective, observer-independent description of quantum dynamics without invoking the collapse postulate, thereby offering a coherent ontological framework for quantum cosmology and a potential mechanism for a non-singular bounce [16, 34].
@@
Building upon this dBB framework, in Ref. [35] a concrete cosmological model featuring a stiff-matter dominated quantum bounce that transitions smoothly into a standard inflationary phase was developed. The model, which considers a flat Friedmann-Lemaître-Robertson-Walker (FLRW) Universe with a scalar field, yields complete analytical solutions spanning the entire history from quantum contraction through the bounce to the end of inflation. This analytical tractability provides a powerful tool for investigating how fundamental quantum parameters of the pre-bounce wave function — such as the characteristic bounce density and wave packet dispersion — determine the number of inflationary e-folds and thereby set the initial conditions for the formation of large-scale structure.
@@ II.2 Quantum Model: de Broglie–Bohm Interpretation
which leads, upon imposing the Hamiltonian constraint, to the Wheeler–DeWitt (WDW) equation in the kination regime,
@@ III Primordial Scalar Perturbations
The primary challenge is to evolve quantum vacuum fluctuations through the non-singular bounce and into the subsequent inflationary epoch. The pre-inflationary dynamics near the bounce imprint a characteristic signature on the perturbation modes that persists throughout the subsequent evolution. To model this process, we segment the history into three distinct dynamical regimes where the background evolution can be well-approximated, thus permitting a tractable analytical treatment for the perturbations:
@@ III.1 Bounce phase
where $\alpha$ and $\mathcal{V}_{0}$ are free parameters fixed in such a way that Eq. (20) can become a good approximation to Eq. (LABEL:VT). Since we can analytically solve the modes equation Eq. (15) for a function like (20), we expect this to provide a good approximation for the solution in the bounce phase.
@@ III.2 Transition phase
During the transition phase, the Universe smoothly transits from an equation of state with $\omega=1$ (kination) to $\omega=-1$ (inflation). Hence, we can consider $k^{2}\gg a^{\prime\prime}(\eta)/a(\eta)$ in Eq. (15) and the solution for the modes $\mu_{k}$ is simply plane-wave like,
@@ III.3 Inflationary phase
where $\alpha_{k}$ and $\beta_{k}$ are arbitrary constants, $H^{\rm(1)}_{\nu}(z)$ and $H^{\rm(2)}_{\nu}(z)$ are the Hankel functions of the first and second kinds, respectively.
@@ III.4 Matching the solutions
Let us now determine all the arbitrary coefficients that appear in the solutions obtained above in each of the three phases. We are in particular interested in the coefficients $\alpha_{k}$ and $\beta_{k}$ of the slow-roll inflationary modes, Eq. (26), resulting from the whole evolution beginning far before the bounce, in the contracting phase. All the coefficients are obtained by matching the solutions and are explained below.
@@
We now match the bounce modes solution in the far past with the one in the next phase, the transition phase. This is done by computing the bounce modes, Eq. (III.1), for $\eta-\eta_{\rm B}\gg 0$. To do this, we first note that in the limit $\eta-\eta_{\rm B}\gg 0$,
@@
Hence, the bounce modes result Eq. (III.1) for $\eta-\eta_{\rm B}\gg 0$ can be simplified and we find
@@ IV Analysis and Results
The constraints obtained for the standard cosmological parameters that exhibit a Gaussian-like posterior profile are presented in Table 1, where we also report the total matter density ($\Omega_{m}$) and the amplitude of matter fluctuations ($\sigma_{8}$) parameters. The corresponding 2D confidence levels (CLs) and 1D marginalized posterior distributions are plotted in Fig. 4.
@@
It should be noted that our cosmological analysis also sheds light on the well-known $H_{0}$-$\sigma_{8}$ correlation and its potential breaking within a QC model. As detailed in Table 1, for the $c=0.003$ case, our results indicate a $H_{0}$ value that remains consistent with those found for the $c=0.03$ and $c=0.3$ scenarios. Simultaneously, we observe a slightly lower $\sigma_{8}$ value compared to the higher $c$ values. This particular shift, exhibiting a tendency towards an anti-correlation between $H_{0}$ and $\sigma_{8}$, holds significant relevance in the context of current cosmological tensions [49].
@@ V Conclusions
In this work, we have investigated the pre-inflationary Universe in the context of quantum cosmology using the de Broglie-Bohm interpretation. Our primary goal was to analyze the evolution of primordial perturbations during a quantum bounce phase and to constrain the model parameters through the observational implications of this theoretical framework.
@@
Finally, we have shown a possible connection of our QC model with the $H_{0}$ and $\sigma_{8}$ tensions. By driving $H_{0}$ towards values more consistent with local measurements, while concurrently achieving a relatively lower $\sigma_{8}$ (compared to other parameter choices within our model), our QC model offers a promising avenue to reconcile these long-standing discrepancies between early- and late-Universe observations of cosmic evolution. This aligns with recent discussions in the literature, where analogous mechanisms in inflationary models have been explored to break such correlations and mitigate cosmological tensions, as discussed for example in Ref. [54].
@@ Acknowledgements.
MB acknowledge the support of Istituto Nazionale di Fisica Nucleare (INFN), iniziativa specifica QGSKY. We thank the use of the CAMB code [36] and the Cobaya MCMC sampler [37] in this work. R.O.R. is partially supported by research grants from Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq), Grant No. 307286/2021-5, and Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ), Grant No. E-26/201.150/2021.